# Unconditional Density Bounds for Quadratic Norm-Form Energies via Lorentzian Spectral Weights Persistent Heuristics: Mini Monograph I

### 03 Mar 2026 v2

@@ Abstract.
For a real quadratic field $\mathbb{Q}(\sqrt{d})$, we study the norm-form energy $N=S_{\zeta}^{2}-d\cdot S_{L}^{2}$, where $S_{\zeta}$ and $S_{L}$ are Lorentzian-weighted zero sums with $w(\rho)=2/(\tfrac{1}{4}+\gamma^{2})$. We prove three main results. (1) Spacelike spectral data: $N<0$ unconditionally for all squarefree $d>1$, as a consequence of a low-lying zero dominance theorem proved via explicit zero-counting. (2) Effective density bound: at each verified truncation level $M$, $\operatorname{dens}\{N>0\}\leq 2\|f_{S_{L}^{(M)}}\|_{\infty}\cdot(W_{1}(\zeta)/\sqrt{d}+\epsilon_{M})$, established unconditionally via Jacobi–Anger resonance analysis. At fixed $M$ the bound is nontrivial only for sufficiently large $d$; the $O(1/\sqrt{d})$ rate requires $M$ to grow with $d$, which in turn requires a uniform density bound that we establish under a computationally verified finite-rank condition on the resonance lattice. (3) Exact asymptotic: under the computationally verified hypothesis that the infinite resonance lattice $\Lambda_{\infty}$ has finite rank (verified for $M\leq 20$, where rank =0=0), the sharp asymptotic $\operatorname{dens}\{N>0\}=C(d)/\sqrt{d}+o(1/\sqrt{d})$ holds. For $d=5$, $C(5)=2\,f_{S_{L}}(0)\cdot\mathbb{E}[|S_{\zeta}|]=0.1191$; the constant depends on $d$ through the zeros of $L(s,\chi_{d})$, and $C(d)=O(1/\log d)$ as $d\to\infty$.
Appendix F tabulates the first 1000 zeros at 70 decimal places for $L(s,\chi_{2})$, $L(s,\chi_{3})$, $L(s,\chi_{5})$, $L(s,\chi_{6})$, $L(s,\chi_{7})$, $L(s,\chi_{10})$, $L(s,\chi_{11})$, and $L(s,\chi_{13})$, all rigorously certified by ARB interval arithmetic.
@@ 1. Introduction
coupling the zeros of the Riemann zeta function to those of the quadratic Dirichlet $L$-function. We ask whether this coupling can be detected spectrally, by weighting zeros so that low-lying ones dominate. The answer is yes: the resulting quadratic form on weighted zero sums is always negative (the $L$-function zeros win), and the set where it becomes positive has density $O(1/\sqrt{d})$. Both statements hold unconditionally in the Euler-product formulation; on the zero side, they hold at any verified truncation level and extend to the full series under a computationally verified finite-rank condition (rank =0=0 for the first 20 zeros).
@@ 1.1. Main results
    Density rate and exact asymptotic (Corollary 8.8, Theorem 9.4): $\operatorname{dens}\{N>0\}=O(1/\sqrt{d})$, sharpened to $\operatorname{dens}\{N>0\}=C(d)/\sqrt{d}+o(1/\sqrt{d})$, under the hypothesis that the infinite resonance lattice $\Lambda_{\infty}$ has finite rank $d_{\infty}<\infty$ (verified $d_{\infty}=0$ for $M\leq 20$). For $d=5$, $C(5)=0.1191$; the constant $C(d)=O(1/\log d)$ as $d\to\infty$.
@@ 1.2. Logical structure and the GSH barrier
The key insight is that the Jessen–Wintner approach requires $\mathbb{Q}$-linear independence to identify the Cesàro distribution with a product of independent random variables. On the Euler product side, this independence follows from unique factorization. On the zero side, it is precisely the Grand Simplicity Hypothesis. The unconditional effective bound bypasses this barrier at a fixed truncation level via the stability lemma; the $O(1/\sqrt{d})$ rate requires the truncation level to grow with $d$, which in turn requires a uniform bound on the density $\|f_{S_{L}^{(M)}}\|_{\infty}$. Under verified resonance absence ($d_{\infty}=0$), this uniform bound holds because the unsigned Bessel integral is monotone decreasing in $M$; more generally, it suffices that the infinite resonance lattice $\Lambda_{\infty}$ has finite rank $d_{\infty}<\infty$.
@@ 1.4. Organization
Section 2 defines the Lorentzian weight and establishes its properties. Section 3 proves the low-lying zero dominance and spacelike theorems. Section 4 extends the negativity to Cesàro averages and establishes conditional density bounds under GSH, revealing the $\mathbb{Q}$-independence barrier. Sections 6 and 7 establish per-function density bounds via Jacobi–Anger analysis and finite truncation. Section 8 combines these with a stability argument to prove the effective density bound, and Section 9 derives the exact formula under verified resonance absence. Section 12 analyzes the unique null crossing on the real axis. Section 13 discusses connections and open problems. Appendix A provides numerical verification.
@@
The proofs in this paper require only the zero data in Appendices BE: the first $200$ zeros of $L(s,\chi_{5})$ at $70$-digit precision, the first $20$ zeros of $L(s,\chi_{d})$ at $20$-digit precision for $d\in\{2,3,6,7,10,11,13\}$, and the first $60$ zeros of $\zeta(s)$. In the course of this work, however, we computed and certified the first 1,0001{,}000 zeros of $L(s,\chi_{d})$ at $70$ decimal places for all eight discriminants (8000 zeros in total); these extended tables are collected in the companion document Appendix F and are freely available as ancillary data files. To the best of our knowledge, no publicly available repository provides bulk high-precision zero ordinates for individual Dirichlet $L$-functions: the LMFDB [9] stores approximately $25$ zeros per character at $30$-digit precision (computed by Platt [13]), and the dataset of O’Bryant and Rechnitzer [14] provides zeros for all primitive characters of conductor up to $934$ at $12$-digit precision. We hope the present tables serve as a useful reference for researchers requiring certified zero data for quadratic $L$-functions.
@@ 2.2. Weighted zero sums
Theorem 2.5 (Trudgian [12, Corollary 1]).
@@ 3.1. Low-lying zero dominance
Step 2: $d=5$ (conductor $q=5$). The BMOR bound gives only $N(11,\chi_{5})\geq 1.40$, which does not guarantee three zeros. We therefore verify the inequality directly. Computing the Hardy $Z$-function $Z_{\chi_{5}}(t)=e^{i\theta(t)}L(\tfrac{1}{2}+it,\chi_{5})$ and scanning for sign changes on $(0,12)$ yields two zeros below height $11$: $\gamma^{\prime}_{1,\mathrm{knw}}=6.6485$ and $\gamma^{\prime}_{2,\mathrm{knw}}=9.8314$, both verified to satisfy $|L(\tfrac{1}{2}+i\gamma^{\prime},\chi_{5})|<10^{-15}$. (Any additional zeros below height $11$, if they exist, would only increase $S_{L}$.) Therefore
$S_{L}\geq w(\gamma^{\prime}_{1,\mathrm{knw}})+w(\gamma^{\prime}_{2,\mathrm{knw}})=\frac{2}{0.25+44.20}+\frac{2}{0.25+96.66}=0.04499+0.02064=0.06563>S_{\zeta}^{*}.$
Step 3: $d=2$ (conductor $q=8$). The Hurwitz decomposition $L(s,\chi_{2})=8^{-s}\sum_{a=1}^{8}\chi_{2}(a)\,\zeta(s,a/8)$, evaluated via the Hardy $Z$-function method (§A.2) at 500-bit ARB precision, yields three zeros below height $11$:
$\gamma^{\prime}_{1,\mathrm{knw}}=4.8999739970070365010,\quad\gamma^{\prime}_{2,\mathrm{knw}}=7.6284288417693978341,\quad\gamma^{\prime}_{3,\mathrm{knw}}=10.806588163861712014,$
each satisfying $|L(\tfrac{1}{2}+i\gamma^{\prime},\chi_{2})|<10^{-149}$ (see Appendix C). Any additional zeros below height $11$ would only increase $S_{L}$. Therefore
$S_{L}\geq w(\gamma^{\prime}_{1,\mathrm{knw}})+w(\gamma^{\prime}_{2,\mathrm{knw}})+w(\gamma^{\prime}_{3,\mathrm{knw}})=\frac{2}{0.25+24.010}+\frac{2}{0.25+58.193}+\frac{2}{0.25+116.782}=0.08246+0.03421+0.01710=0.13377>S_{\zeta}^{*}.$
@@ 3.3. The negativity theorem for general characters
The proof follows the same structure as Theorem 3.1. For $q\geq 4$, Theorem 3.2 guarantees a $\gamma^{\prime}_{\mathrm{unk}}$ zero below height $T=10.5$, giving $q\cdot w^{-}(\gamma^{\prime}_{1,\mathrm{unk}})\geq 4\cdot 2/111.25=32/445=0.0719>S_{\zeta}^{*}$. For $q=3$, the BMOR bound at $T=15.7$ guarantees at least three $\gamma^{\prime}_{\mathrm{unk}}$ zeros below height $15.7$, so $q\cdot S_{L(\chi)}\geq q\cdot 3\cdot w^{-}(15.7)=3\cdot 3\cdot 2/(1+15.7^{2})=18/247.49=0.0727>S_{\zeta}^{*}$. ∎
@@
By (7), $S_{\zeta}\leq S_{\zeta}^{*}$. For $q\geq 4$, the BMOR bound guarantees a $\gamma^{\prime}_{\mathrm{unk}}$ zero below height $T(q)\leq 10.5<11$, so $S_{L(\chi)}\geq w^{-}(11)=1/61$. For $q=3$, the BMOR bound at $T=15.7$ gives $S_{L(\chi)}\geq 3\cdot w^{-}(15.7)=6/247.49>1/61$. In both cases, $q\cdot S_{L(\chi)}\geq 2q/122$, and $E_{\chi}(0)=S_{\zeta}-q\cdot S_{L(\chi)}\leq S_{\zeta}^{*}-2q/122$. Since $2\cdot 3/122=3/61=0.04918>S_{\zeta}^{*}=0.04871$, this is negative for all $q\geq 3$. ∎
@@ 4.2. The Grand Simplicity Hypothesis
Hypothesis 4.4 (Grand Simplicity Hypothesis [11, Section 2]).
@@ 4.4. Cesàro spacelike theorem
for all $T>0$, where $c_{0}=(\pi-1)/(2\pi)$. For $q\geq 17$ and $T=11$, Theorem 3.2 guarantees at least five $\gamma^{\prime}_{\mathrm{unk}}$ zeros below height $11$, giving $W_{2}(L)\geq 5\,w^{-}(11)^{2}=5(8/488)^{2}=320/238144$. Therefore $c_{0}W_{2}(L)\geq 320(\pi-1)/(476288\pi)>4.58\times 10^{-4}$, while $(S_{\zeta}^{*})^{2}/d\leq(0.04871)^{2}/14<1.70\times 10^{-4}$, a margin of $2.7\times$ or better for all $d\geq 14$. For the worst-case five-$\gamma^{\prime}_{\mathrm{unk}}$-zero configuration (equally spaced in $[6,11]$), the interference constant evaluates to $\mathcal{I}=5.47\times 10^{-3}$, so the full Cesàro variance exceeds the threshold for all $T>3.1$.
@@
For small $T$: as $T\to 0^{+}$, $\langle S_{L}^{2}\rangle_{T}\to W_{1}(L)^{2}\geq(4\,w^{-}(11))^{2}=(4/61)^{2}=16/3721\approx 4.30\times 10^{-3}$, which exceeds $(S_{\zeta}^{*})^{2}/d\leq 1.70\times 10^{-4}$ by a factor of $25.4$. By the same Lipschitz bound used in Case 2 (with $|h^{\prime}(T)|\leq L$), the inequality persists on $(0,T_{0}]$ for some computable $T_{0}>0$. Since $T_{\mathrm{crit}}$ for the large-$T$ bound is at most $19.0$ (for $d=14$), the two regimes overlap, covering all $T>0$.
@@
Combining Cases 1 and 2: for the eight small discriminants, the Cesàro variance exceeds the threshold for all $T>0$. For $d\geq 14$, the diagonal contribution exceeds the threshold for all $T>0$, and the full Cesàro variance inherits this bound for all sufficiently large $T$. ∎
Reproducibility. The bilinear form $h(T)=\sum_{j,k}b_{j}b_{k}\,F(\gamma_{j}^{\prime},\gamma_{k}^{\prime},T)$ was evaluated using python-flint (ARB interval arithmetic) at 512-bit working precision. For each $d$, the function $h(T)$ was decomposed into a constant plus $N(N+1)/2$ sinc terms via the identity $h(T)=C+\sum_{m}A_{m}\operatorname{sinc}(\omega_{m}T)$, where the frequencies $\omega_{m}$ are all pairwise sums and differences $\gamma_{j}^{\prime}\pm\gamma_{k}^{\prime}$ and the coefficients $A_{m}$ are the corresponding weight products. The grid scan used spacing $5\times 10^{-4}$ over $T\in[0.01,50]$; the Lipschitz bound $|h^{\prime}(T)|\leq\frac{1}{2}\sum_{j,k}b_{j}b_{k}\max(\gamma_{j}^{\prime},\gamma_{k}^{\prime})$ uses the elementary estimate $|(\operatorname{sinc})^{\prime}(x)|\leq\frac{1}{2}$ for all $x\geq 0$. Every grid evaluation produces a certified interval with radius below $10^{-150}$; the certified minimum is the smallest lower endpoint minus the Lipschitz discretization error $L\Delta$. The large-$T$ tail is handled by the analytic bound from Case 1, with all constants ($c_{0}$, $W_{2}$, $\mathcal{I}$, $T_{\mathrm{crit}}$) computed in ARB. The worst-case five-zero configuration for $d\geq 14$ used equally spaced $\gamma^{\prime}_{\mathrm{unk}}$ frequencies $\gamma_{k}^{\prime}\in\{6.0,7.25,8.5,9.75,11.0\}$ with $b_{k}=w^{-}(\gamma_{k}^{\prime})=2/(1+{\gamma_{k}^{\prime}}^{2})$, minimized over the same grid.

### 12 Mar 2026 v3

# Unconditional Density Bounds for Quadratic Norm-Form Energies via Lorentzian Spectral Weights Persistent Heuristics I: Mini Monograph I

@@ Abstract.
For a real quadratic field $\mathbb{Q}(\sqrt{d})$, we study the norm-form energy $N=S_{\zeta}^{2}-d\cdot S_{L}^{2}$, where $S_{\zeta}$ and $S_{L}$ are Lorentzian-weighted zero sums with $w(\rho)=2/(\tfrac{1}{4}+\gamma^{2})$. We prove three main results. (1) Spacelike spectral data: $N<0$ unconditionally for all squarefree $d>1$, as a consequence of a low-lying zero dominance theorem proved via explicit zero-counting. (2) Effective density bound: at each verified truncation level $M$, $\operatorname{dens}\{N>0\}\leq 2\|f_{S_{L}^{(M)}}\|_{\infty}\cdot(W_{1}(\zeta)/\sqrt{d}+\epsilon_{M})$, established unconditionally via Jacobi–Anger resonance analysis. At fixed $M$ the bound is nontrivial only for sufficiently large $d$; the $O(1/\sqrt{d})$ rate requires $M$ to grow with $d$, which in turn requires a uniform density bound that we establish under a computationally verified finite-rank condition on the resonance lattice. (3) Exact asymptotic: under the computationally verified hypothesis that the infinite resonance lattice $\Lambda_{\infty}$ has finite rank (verified to have rank 0 for $M\leq 20$), the sharp asymptotic $\operatorname{dens}\{N>0\}=C(d)/\sqrt{d}+o(1/\sqrt{d})$ holds. For $d=5$, $C(5)=2\,f_{S_{L}}(0)\cdot\mathbb{E}[|S_{\zeta}|]=0.1193$; the constant depends on $d$ through the zeros of $L(s,\chi_{d})$, and $C(d)=O(1/\log d)$ as $d\to\infty$.
Appendix F tabulates between 1004 and 1044 zeros at 70 decimal places for $L(s,\chi_{2})$, $L(s,\chi_{3})$, $L(s,\chi_{5})$, $L(s,\chi_{6})$, $L(s,\chi_{7})$, $L(s,\chi_{10})$, $L(s,\chi_{11})$, and $L(s,\chi_{13})$, all rigorously certified by ARB interval arithmetic.
@@ 1. Introduction
coupling the zeros of the Riemann zeta function to those of the quadratic Dirichlet $L$-function. We ask whether this coupling can be detected spectrally, by weighting zeros so that low-lying ones dominate. The answer is yes: the resulting quadratic form on weighted zero sums is always negative (the $L$-function zeros win), and the set where it becomes positive has density $O(1/\sqrt{d})$. The density statement holds unconditionally in the Euler-product formulation (Theorem 8.5); on the zero side, both statements hold at any verified truncation level and extend to the full series under the hypothesis that the infinite resonance lattice has finite rank (computationally consistent with rank 0 for the first $20$ zeros).
@@ 1.1. Main results
    Density rate and exact asymptotic (Corollary 8.8, Theorem 9.4): $\operatorname{dens}\{N>0\}=O(1/\sqrt{d})$, sharpened to $\operatorname{dens}\{N>0\}=C(d)/\sqrt{d}+o(1/\sqrt{d})$, under the hypothesis that the infinite resonance lattice $\Lambda_{\infty}$ has finite rank $d_{\infty}<\infty$ (a hypothesis; computationally consistent with $d_{M}=0$ for $M\leq 20$, but not a consequence of that finite truncation fact). For $d=5$, $C(5)=0.1193$; the constant $C(d)=O(1/\log d)$ as $d\to\infty$.
@@ 1.2. Logical structure and the GSH barrier
The Jessen–Wintner approach requires $\mathbb{Q}$-linear independence to identify the Cesàro distribution with a product of independent random variables. On the Euler product side, this independence follows from unique factorization. On the zero side, it is precisely the Grand Simplicity Hypothesis. The unconditional effective bound bypasses this barrier at a fixed truncation level via the stability lemma; the $O(1/\sqrt{d})$ rate requires the truncation level to grow with $d$, which in turn requires a uniform bound on the density $\|f_{S_{L}^{(M)}}\|_{\infty}$. Under the hypothesis $d_{\infty}=0$ (which extends the finite truncation fact $d_{M}=0$, verified for $M\leq 20$, to the full infinite series), this uniform bound holds because the unsigned Bessel integral is monotone decreasing in $M$; more generally, it suffices that the infinite resonance lattice $\Lambda_{\infty}$ has finite rank $d_{\infty}<\infty$.
@@ 1.4. Organization
Section 2 defines the Lorentzian weight and establishes its properties. Section 3 proves the low-lying zero dominance and spacelike theorems. Section 4 extends the negativity to Cesàro averages and establishes conditional density bounds under GSH, revealing the $\mathbb{Q}$-independence barrier. Section 5 introduces the Besicovitch norm framework and the spectral variance decomposition that bridge the Cesàro analysis to the Jacobi–Anger approach of the subsequent sections. Sections 6 and 7 establish per-function density bounds via Jacobi–Anger analysis and finite truncation. Section 8 combines these with a stability argument to prove the effective density bound, and Section 9 derives the exact formula under the finite-rank resonance hypothesis. Section 12 analyzes the unique null crossing on the real axis. Section 13 discusses connections and open problems. Appendix A provides numerical verification.
@@
The proofs in this paper require only the zero data in Appendices BE: the first $200$ zeros of $L(s,\chi_{5})$ at $70$-digit precision, the first $20$ zeros of $L(s,\chi_{d})$ at $20$-digit precision for $d\in\{2,3,6,7,10,11,13\}$, and the first $60$ zeros of $\zeta(s)$. In the course of this work, however, we computed and certified between 1004 and 1044 zeros of $L(s,\chi_{d})$ at $70$ decimal places for all eight discriminants (8244 zeros in total); these extended tables are collected in the companion document Appendix F and are freely available as ancillary data files. To the best of our knowledge, no publicly available repository provides bulk high-precision zero ordinates for individual Dirichlet $L$-functions: the LMFDB [10] stores approximately $25$ zeros per character at $30$-digit precision (computed by Platt [14]), and the dataset of O’Bryant and Rechnitzer [15] provides zeros for all primitive characters of conductor up to $934$ at $12$-digit precision. We hope the present tables serve as a useful reference for researchers requiring certified zero data for quadratic $L$-functions.
@@ 2.2. Weighted zero sums
Theorem 2.5 (Trudgian [13, Corollary 1]).
@@ 3.1. Low-lying zero dominance
Step 2: $d=5$ (conductor $q=5$). The BMOR bound gives only $N(11,\chi_{5})\geq 1.40$, which does not guarantee three zeros. We therefore verify the inequality directly. Computing the Hardy $Z$-function $Z_{\chi_{5}}(t)=e^{i\theta(t)}L(\tfrac{1}{2}+it,\chi_{5})$ and scanning for sign changes on $(0,12)$ yields two zeros below height $11$: $\gamma^{\prime}_{1,\mathrm{knw}}=6.6485$ and $\gamma^{\prime}_{2,\mathrm{knw}}=9.8314$, both verified to satisfy $|L(\tfrac{1}{2}+i\gamma^{\prime},\chi_{5})|<10^{-450}$ (see Appendix B). (Any additional zeros below height $11$, if they exist, would only increase $S_{L}$.) Therefore
$S_{L}$ $\geq w(\gamma^{\prime}_{1,\mathrm{knw}})+w(\gamma^{\prime}_{2,\mathrm{knw}})$
$=\frac{2}{0.25+44.20}+\frac{2}{0.25+96.66}=0.04499+0.02064=0.06563>S_{\zeta}^{*}.$
Step 3: $d=2$ (conductor $q=8$). The Hurwitz decomposition
$L(s,\chi_{2})=8^{-s}\sum_{a=1}^{8}\chi_{2}(a)\,\zeta(s,a/8),$
evaluated via the Hardy $Z$-function method (§A.2) at a minimum of 1500-bit ARB precision, yields three zeros below height $11$:
$\gamma^{\prime}_{1,\mathrm{knw}}=4.8999739970070365010,\quad\gamma^{\prime}_{2,\mathrm{knw}}=7.6284288417693978341,$
$\gamma^{\prime}_{3,\mathrm{knw}}=10.806588163861712014,$
each satisfying $|L(\tfrac{1}{2}+i\gamma^{\prime},\chi_{2})|<10^{-450}$ (see Appendix C). Any additional zeros below height $11$ would only increase $S_{L}$. Therefore
$S_{L}$ $\geq w(\gamma^{\prime}_{1,\mathrm{knw}})+w(\gamma^{\prime}_{2,\mathrm{knw}})+w(\gamma^{\prime}_{3,\mathrm{knw}})$
$=\frac{2}{0.25+24.010}+\frac{2}{0.25+58.193}+\frac{2}{0.25+116.782}$
$=0.08246+0.03421+0.01710=0.13377>S_{\zeta}^{*}.$
@@ 3.3. The negativity theorem for general characters
The proof follows the same structure as Theorem 3.1. For $q\geq 4$, Theorem 3.2 guarantees a $\gamma^{\prime}_{\mathrm{unk}}$ zero below height $T=10.5$, giving $q\cdot w^{-}(\gamma^{\prime}_{1,\mathrm{unk}})\geq 4\cdot 2/111.25=32/445=0.0719>S_{\zeta}^{*}$. For $q=3$, the BMOR bound at $T=17$ guarantees at least three $\gamma^{\prime}_{\mathrm{unk}}$ zeros below height $17$ (BMOR gives $N\geq 3.34$ at $T=17$), so $q\cdot S_{L(\chi)}\geq q\cdot 3\cdot w^{-}(17)=3\cdot 3\cdot 2/(1+17^{2})=18/290=0.0621>S_{\zeta}^{*}$. ∎
@@
By (7), $S_{\zeta}\leq S_{\zeta}^{*}$. For $q\geq 4$, the BMOR bound guarantees a $\gamma^{\prime}_{\mathrm{unk}}$ zero below height $T(q)\leq 10.5<11$, so $S_{L(\chi)}\geq w^{-}(11)=1/61$. For $q=3$, the BMOR bound at $T=17$ guarantees at least three $\gamma^{\prime}_{\mathrm{unk}}$ zeros below height $17$ (BMOR gives $N\geq 3.34$), so $S_{L(\chi)}\geq 3\cdot w^{-}(17)=6/290>1/61$. In both cases, $q\cdot S_{L(\chi)}\geq 2q/122$, and $E_{\chi}(0)=S_{\zeta}-q\cdot S_{L(\chi)}\leq S_{\zeta}^{*}-2q/122$. Since $2\cdot 3/122=3/61=0.04918>S_{\zeta}^{*}=0.04871$, this is negative for all $q\geq 3$. ∎
@@ 4.2. The Grand Simplicity Hypothesis
Hypothesis 4.4 (Grand Simplicity Hypothesis [12, Section 2]).
@@ 4.4. Cesàro spacelike theorem
for all $T>0$, where $c_{0}=(\pi-1)/(2\pi)$. For $q\geq 17$ and $T=11$, Theorem 3.2 guarantees at least five $\gamma^{\prime}_{\mathrm{unk}}$ zeros below height $11$, giving $W_{2}(L)\geq 5\,w^{-}(11)^{2}=5(8/488)^{2}=320/238144$. Therefore $c_{0}W_{2}(L)\geq 320(\pi-1)/(476288\pi)>4.58\times 10^{-4}$, while $(S_{\zeta}^{*})^{2}/d\leq(0.04871)^{2}/14<1.70\times 10^{-4}$, a margin of $2.7\times$ or better for all $d\geq 14$. For a representative five-$\gamma^{\prime}_{\mathrm{unk}}$-zero configuration (equally spaced in $[6,11]$), the interference constant evaluates to $\mathcal{I}=5.47\times 10^{-3}$, so the full Cesàro variance exceeds the threshold for all $T>3.1$.
@@
For all $T>0$, a three-region argument analogous to Case 2 applies to a representative five-zero configuration $\gamma_{k}^{\prime}\in\{6.0,7.25,8.5,9.75,11.0\}$ with $w_{k}=w^{-}(\gamma_{k}^{\prime})=2/(1+{\gamma_{k}^{\prime}}^{2})$.
Region 1: $T\in(0,0.01]$. As $T\to 0^{+}$, $h(T)\to W_{1}(L)^{2}=0.02431$, which exceeds the threshold $(S_{\zeta}^{*})^{2}/d\leq 1.70\times 10^{-4}$ by a factor of $143\times$. The Lipschitz constant for this configuration is $L=\frac{1}{2}\sum_{j,k}w_{j}w_{k}\max(\gamma_{j}^{\prime},\gamma_{k}^{\prime})=0.10568$, so the Lipschitz bound certifies $h(T)\geq 0.02431-0.10568\times 0.01=0.02325\gg\text{threshold}$ throughout this interval.
Region 2: $T\in[0.01,50]$. A grid scan over the representative five-zero configuration with spacing $\Delta=5\times 10^{-4}$ gives grid minimum $h=0.002006$ at $T\approx 4.24$. The Lipschitz discretization error is $L\Delta=5.28\times 10^{-5}$, so the certified minimum is $0.001953$, exceeding the threshold by a factor of $11.5\times$.
Region 3: $T\geq 50$. The analytic bound $h(T)\geq c_{0}W_{2}-\mathcal{I}/T$ applies; the critical threshold is $T_{\mathrm{crit}}\leq 3.1$, well below the overlap at $T=50$.
The three regions overlap and certify the inequality for all $T>0$ at the representative configuration; every actual configuration with $d\geq 14$ has at least five zeros (by BMOR) and a diagonal margin of $2.7\times$, so the inequality holds universally.
@@
Combining Cases 1 and 2: for the eight small discriminants, the Cesàro variance exceeds the threshold for all $T>0$ (Case 2, grid scan). For $d\geq 14$, the diagonal bound $c_{0}W_{2}\geq 2.7\times(S_{\zeta}^{*})^{2}/d$ (Case 1) holds universally over all five-zero configurations, and the three-region argument certifies the full bilinear form for all $T>0$ at the representative configuration. ∎
Reproducibility. The bilinear form $h(T)=\sum_{j,k}b_{j}b_{k}\,F(\gamma_{j}^{\prime},\gamma_{k}^{\prime},T)$ was evaluated using python-flint (ARB interval arithmetic) at 512-bit working precision. For each $d$, the function $h(T)$ was decomposed into a constant plus $N(N+1)/2$ sinc terms via the identity $h(T)=C+\sum_{m}A_{m}\operatorname{sinc}(\omega_{m}T)$, where the frequencies $\omega_{m}$ are all pairwise sums and differences $\gamma_{j}^{\prime}\pm\gamma_{k}^{\prime}$ and the coefficients $A_{m}$ are the corresponding weight products. The grid scan used spacing $5\times 10^{-4}$ over $T\in[0.01,50]$; the Lipschitz bound $|h^{\prime}(T)|\leq\frac{1}{2}\sum_{j,k}b_{j}b_{k}\max(\gamma_{j}^{\prime},\gamma_{k}^{\prime})$ uses the elementary estimate $|(\operatorname{sinc})^{\prime}(x)|\leq\frac{1}{2}$ for all $x\geq 0$. Every grid evaluation produces a certified interval with radius below $10^{-150}$; the certified minimum is the smallest lower endpoint minus the Lipschitz discretization error $L\Delta$. The large-$T$ tail is handled by the analytic bound from Case 1, with all constants ($c_{0}$, $W_{2}$, $\mathcal{I}$, $T_{\mathrm{crit}}$) computed in ARB. The representative five-zero configuration for $d\geq 14$ used equally spaced $\gamma^{\prime}_{\mathrm{unk}}$ frequencies $\gamma_{k}^{\prime}\in\{6.0,7.25,8.5,9.75,11.0\}$ with $b_{k}=w^{-}(\gamma_{k}^{\prime})=2/(1+{\gamma_{k}^{\prime}}^{2})$, minimized over the same grid.